A passenger train leave the station 4 hours after the freight train left the same station. The freight train is traveling 35km/h slower than the passenger train. Find the speed of passenger train if it over takes the freight train in five hours.
Answer by ankor@dixie-net.com(22740)

passenger train leave the station 4 hours after the freight train left the same station.
The freight train is traveling 35km/h slower than the passenger train.
Find the speed of passenger train if it over takes the freight train in five hours.
:
When the pass train passes the freight, they will have gone the same distance.
From the information given, we know that the pass train traveled 5 hr and the freight traveled 9 hrs
:
let s = the speed of the pass train
then
(s-35) = speed of the freight
;
Write a dist equation, dist = time * speed
5s = 9(s-35)
5s = 9s - 315
5s - 9s = -315
-4s = -315
s = -315/-4
s = 78.75 mph is the speed of the pass train
:
:
Check; find the dist of each train, should be equal
Freight speed: 78.75 - 35 = 43.75
5 * 78.75 = 393.75
9 * 43.75 = 393.75